OpenStudy (anonymous):

Solve the system of equations by graphing. Then classify the system. x+y=11 x-y=-7

OpenStudy (anonymous):

The solution for this question is non. so I need help?

OpenStudy (anonymous):

well easiest way is making everythign in terms of y. (or x wtv is easier) so you get y=-x+11 and y=x+7 then you just plug in two points for each to graph it because that is all you need when the lines are linear.

OpenStudy (anonymous):

then you compare the two so you get x+7=-x+11 so your x=2 and then plug into y and you get 9 get it :)?

OpenStudy (anonymous):

okay, may you explain more.

OpenStudy (anonymous):

sure what dont u get?

OpenStudy (anonymous):

Everything becoz I try to solve it but there is any solution? hle plz

OpenStudy (anonymous):

What is the solution for the system?

OpenStudy (anonymous):

i wrote it x=2 y=9